OpenStudy (anonymous):

Determine an equation of the line through the given points. (-3, 5), (1,7).

OpenStudy (anonymous):

(Y2-Y1)/(X2-X1)=x i believe

OpenStudy (anonymous):

or =m then its y=mx+b

OpenStudy (anonymous):

I am sorry, i'm not getting it.

OpenStudy (anonymous):

I got m=1/2.........i don't know where to go after that.

OpenStudy (anonymous):

This book is Terrible

OpenStudy (anonymous):

Slope is rise over run. Rise = change in the y-coordinate, run=change in the x-coordinate. You have two points. Call them (x1,y1), (x2,y2). So you plug them into (y1-y2)/(x1-x2). It doesn't matter what one you do first, as long as both points are consistent. So you got the slope right, so m is 1/2. :) Then you take a point and plug it in to solve for when x=0. Take (-3,5) and plug it in to y=mx+b 5 = .5 * -3 + b 5 = -1.5 + b 6.5 = b y = .5x + 6.5, if I'm right.

OpenStudy (anonymous):

did you try youtubing it?

OpenStudy (anonymous):

very good, now plug in the calculator and see if its correct :D

OpenStudy (anonymous):

Well, I wil YouTube it in a second...thanks..I got 5=1/2x+3